April, Bill , Candace, and Bobby are to be seated at random in a row of 5 chairs. What is the probability that April and Bobby will occupy the seats at the end of the row?
1
Expert's answer
2020-04-30T18:03:30-0400

There are 5 chairs and 4 persons.

Total number of ways(N) the can be sit is,

"N=^5p\ _4=\frac{5!}{1!}=120"


Ways(n) of April and Bobby will occupy the seats at the end of the row is,

"n=2*^3p\ _2=2*6=12"


Therefore,the probability that April and Bobby will occupy the seats at the end of the row,

"P=\frac{n}{N}=\frac{12}{120}=0.1"


Answer is 0.1